# Pointing the Way: Refining Radar-Lidar Localization Using Learned ICP Weights

### v3

@@ Abstract
This paper presents a novel deep-learning-based approach to improve localizing radar measurements against lidar maps. This radar-lidar localization leverages the benefits of both sensors; radar is resilient against adverse weather, while lidar produces high-quality maps in clear conditions. However, owing in part to the unique artefacts present in radar measurements, radar-lidar localization has struggled to achieve comparable performance to lidar-lidar systems, preventing it from being viable for autonomous driving. This work builds on ICP-based radar-lidar localization by including a learned preprocessing step that weights radar points based on high-level scan information. To train the weight-generating network, we present a novel, stand-alone, open-source differentiable ICP library. The learned weights facilitate ICP by filtering out harmful radar points related to artefacts, noise, and even vehicles on the road. Combining an analytical approach with a learned weight reduces overall localization errors and improves convergence in radar-lidar ICP results run on real-world autonomous driving data. Our code base is publicly available to facilitate reproducibility and extensions222 Paper code is available at: https://github.com/utiasASRL/mm_masking..
@@ II RELATED WORK
Radar has seen a resurgence within robotics in recent years [radar_survey, radar_survey_2, venon2022millimeter]. The ability of modern-day radar systems to produce reasonable long-term odometry results is shown in [Cen_Newman_2018], where an ICP-style pointcloud alignment is used to compute the relative motion between consecutive timestamps. An alternative approach is presented in [masking_by_moving], where a network is trained to mask raw radar scans before using them for correlative scan matching of consecutive frames. It is shown that such a learned approach is capable of outperforming heuristic filtering methods. A network that directly extracts features, descriptors, and feature scores from raw radar images and then uses them to match radar data from consecutive timesteps is presented in [under_the_radar]. A self-supervised feature learning approach, which does not require groundtruth data, is shown in [hero_paper]. Although feature-extraction methods are typically more interpretable and robust to poor initial guesses, they frequently require large networks with higher computational cost. Additionally, feature matching across sensing modalities, such as between radar and lidar, poses an extra challenge. Finally, most radar algorithms have focused on odometry or simultaneous localization and mapping (SLAM), e.g., [Jose_Adams_2004, Rouveure_Monod_Faure_2009, Checchin_Gerossier_Blanc_Chapuis_Trassoudaine_2010, Schuster_Keller_Rapp_Haueis_Curio_2016], meaning that investigations into the performance of radar in localization against a previously known map is understudied. An extended Kalman filter (EKF) ICP-based approach for radar-radar localization is presented in [Ward_Folkesson_2016]. However, their localization errors are on the order of meters, far above autonomous driving requirements. In our recent ICP-based approach [are_we_ready_for], localization is performed using batch optimization with a modified Constant False Alarm Rate (CFAR) [cfar] point extractor called Bounded False Alarm Rate (BFAR) [bfar]. This work shows that radar localization against an existing radar map can approach the level needed for autonomous driving, with translation errors of $6-10\,$\mathrm{cm}$$.
@@ III-C Differentiable ICP
A key part of the pipeline is the differentiable ICP (dICP) algorithm, which is needed to backpropagate directly from the ICP error. This algorithm is made publicly available333 dICP code is available at: https://github.com/utiasASRL/dICP and adds to the recent corpus of open-source, differentiable versions of classic robotic algorithms [pypose, grad_slam, binbin_paper]. Indeed, all three of these libraries include a differentiable ICP implementation. However, they are all restrictive in that they only implement either point-to-point or point-to-plane cost terms, do not make use of any robust cost function or only allow for one type, do not accept weighted pointclouds, and, for [grad_slam, binbin_paper], are embedded within a SLAM pipeline, without a clear way to run ICP in a stand-alone fashion.

### v4

@@ Abstract
This paper presents a novel deep-learning-based approach to improve localizing radar measurements against lidar maps. This radar-lidar localization leverages the benefits of both sensors; radar is resilient against adverse weather, while lidar produces high-quality maps in clear conditions. However, owing in part to the unique artefacts present in radar measurements, radar-lidar localization has struggled to achieve comparable performance to lidar-lidar systems, preventing it from being viable for autonomous driving. This work builds on ICP-based radar-lidar localization by including a learned preprocessing step that weights radar points based on high-level scan information. To train the weight-generating network, we present a novel, stand-alone, open-source differentiable ICP library. The learned weights facilitate ICP by filtering out harmful radar points related to artefacts, noise and even vehicles on the road. Combining an analytical approach with a learned weight reduces overall localization errors and improves convergence in radar-lidar ICP results run on real-world autonomous driving data. Our code base is publicly available to facilitate reproducibility and extensions111 Paper code is available at: https://github.com/utiasASRL/mm_masking..
Index Terms:
Autonomous-vehicle localization, radar perception, learned pointcloud filtering.
@@ II RELATED WORK
Radar has seen a resurgence within robotics in recent years [radar_survey, radar_survey_2, venon2022millimeter]. The ability of modern-day radar systems to produce reasonable long-term odometry results is shown in [Cen_Newman_2018], where an ICP-style pointcloud alignment is used to compute the relative motion between consecutive timestamps. An alternative approach is presented in [masking_by_moving], where a network is trained to mask raw radar scans before using them for correlative scan matching of consecutive frames. It is shown that such a learned approach is capable of outperforming heuristic filtering methods. A network that directly extracts features, descriptors, and feature scores from raw radar images and then uses them to match radar data from consecutive timesteps is presented in [under_the_radar]. A self-supervised feature learning approach, which does not require groundtruth data, is shown in [hero_paper]. Although feature-extraction methods are typically more interpretable and robust to poor initial guesses, they frequently require large networks with higher computational cost. Additionally, feature matching across sensing modalities, such as between radar and lidar, poses an extra challenge. Finally, most radar algorithms have focused on odometry or simultaneous localization and mapping (SLAM), e.g., [Jose_Adams_2004, Rouveure_Monod_Faure_2009, Checchin_Gerossier_Blanc_Chapuis_Trassoudaine_2010, Schuster_Keller_Rapp_Haueis_Curio_2016], meaning that investigations into the performance of radar in localization against a previously known map is understudied. An extended Kalman filter (EKF) ICP-based approach for radar-radar localization is presented in [Ward_Folkesson_2016]. However, their localization errors are on the order of meters, far above autonomous driving requirements. In our recent ICP-based approach [are_we_ready_for], localization is performed using batch optimization with a modified Constant False Alarm Rate (CFAR) [cfar] point extractor called Bounded False Alarm Rate (BFAR) [bfar]. This work shows that radar localization against an existing radar map can approach the level needed for autonomous driving [reid2019localization], with translation errors of $6-10\,$\mathrm{cm}$$.
@@ III-C Differentiable ICP
A key part of the pipeline is the differentiable ICP (dICP) algorithm, which is needed to backpropagate directly from the ICP error. This algorithm is made publicly available222 dICP code is available at: https://github.com/utiasASRL/dICP and adds to the recent corpus of open-source, differentiable versions of classic robotic algorithms [pypose, grad_slam, binbin_paper]. Indeed, all three of these libraries include a differentiable ICP implementation. However, they are all restrictive in that they only implement either point-to-point or point-to-plane cost terms, do not make use of any robust cost function or only allow for one type, do not accept weighted pointclouds, and, for [grad_slam, binbin_paper], are embedded within a SLAM pipeline, without a clear way to run ICP in a stand-alone fashion.